# Entropy-based random models for hypergraphs

### v2

@@ Abstract
Let us start by verifying the correctness of the estimations of the number of empty and parallel hyperedges provided by our benchmarks: to this aim, we have considered the values $N=300$ and $L=1000$ (for other values, see Appendix A and Appendix B).
IV.1.1 1. The Random Hypergraph Model
@@
Let us start focusing on our homogeneous benchmark. The left panel of Fig. 2 depicts the trend of $p_{1}=1-\left[1-\left(1-\frac{T}{NL}\right)^{N}\right]^{L}$. The left panel of Fig. 3 depicts the trend of $\frac{\langle N_{0}\rangle}{L}=\left(1-\frac{T}{NL}\right)^{N}=p_{0}$ and $P(d_{\alpha\beta}=0)=\left[1-\frac{2T}{NL}\left(1-\frac{T}{NL}\right)\right]^{N}$. Each quantity has been plotted as a function of $T/NL\in[10^{-6},1]$. The dense (sparse) regime is recovered for large (small) values of $T/NL$. Each dot of Figs. 2 and 3 represents an average taken over an ensemble of $1000$ configurations explicitly sampled from the RHM and is accompanied by the corresponding $95\%$ confidence interval. A first, general comment concerns the agreement between the analytical trends (solid lines) and the numerical estimations (dots), confirming the correctness of our formulas.
The filling threshold.
@@
Although the value of the filling threshold has been determined by inspecting the asymptotic behaviour of $\langle N_{0}\rangle$, the quantity showing the neatest transition from the sparse to the dense regime is $p_{1}\xrightarrow[]{N\to+\infty}1-(1-e^{-h})^{L}$; evaluating it in correspondence of $h_{f}^{\text{RHM}}=\ln L\simeq 6.907$ returns the value $1-(1-1/L)^{L}\simeq 0.6323$ interestingly enough, letting $L$ grow asymptotically as well, the previous value tends to $1-e^{-1}\simeq 0.6321$.
@@
Upon considering that $p_{0}\xrightarrow[]{N\to+\infty}e^{-h}$, evaluating it in correspondence of $h_{f}^{\text{RHM}}=\ln L\simeq 6.907$ returns the value $0.0011/L=$ 0.001: in words, although the filling threshold ensures that each, single hyperedge is empty with an overall small probability, the likelihood of observing at least one, empty hyperedge is still large (i.e. 2/3$\simeq$ 2/3)≃ 2 / 3 ). The steepness of the trend of $p_{1}$, however, suggests it to quickly vanish as the density of 1s in the incidence matrix crosses the value $p_{f}^{\text{RHM}}=h_{f}^{\text{RHM}}/N=\ln L/N\simeq 0.023$.
Figure 3: Left panel: trends of $p_{0}\xrightarrow[]{N\to+\infty}e^{-h}$ (aquamarine line) and $P(d_{\alpha\beta}=0)\xrightarrow[]{N\to+\infty}e^{-2h}$ (dark-blue line) as a function of $T/NL$; evaluating them in correspondence of $h_{r}^{\text{RHM}}=1/2$ (the related, critical density value is indicated in light-blu) returns the values $e^{-{h_{r}^{\text{RHM}}}}=1/\sqrt{e}\simeq 0.606$ and $e^{-{2h_{r}^{\text{RHM}}}}=e^{-1}\simeq 0.367$; evaluating them in correspondence of $h_{f}^{\text{RHM}}=\ln L\simeq 6.907$ (the related, critical density value is indicated in magenta) returns the values $e^{-{h_{f}^{\text{RHM}}}}=1/L=0.001$ and $e^{-{2h_{f}^{\text{RHM}}}}=1/L^{2}=10^{-6}$. The red line indicates the density value corresponding to $h=1$, ensuring that each, expected hyperdegree amounts to 1. Right panel: trends of $p_{0}$, as defined by Eq. (26), and $P_{0}$, i.e. the percentage of pairs of parallel hyperedges. As already noticed, the shift on the right of the trend induced by the HCM is probably due to the presence of small fitnesses, decreasing the magnitude of the set of probability coefficients. The dense (sparse) regime is recovered for large (small) values of $T/NL$. Each dot represents an average taken over an ensemble of $1000$ configurations (explicitly sampled from either the RHM or the HCM) and is accompanied by the corresponding $95\%$ confidence interval. The yellow line indicates the density value in correspondence of which the minimum, expected hyperdegree amounts to 1.
The resolution threshold.
The value of the resolution threshold has been determined by inspecting the asymptotic behaviour of $\langle d_{\alpha\beta}\rangle$: evaluating $P(d_{\alpha\beta}=0)\xrightarrow[]{N\to+\infty}e^{-2h}$ in correspondence of $h_{r}^{\text{RHM}}=1/2$ returns the value $e^{-1}\simeq 0.367$.
@@
Upon considering that $p_{0}\xrightarrow[]{N\to+\infty}e^{-h}$, evaluating it in correspondence of $h_{r}^{\text{RHM}}=1/2$ returns the value $1/\sqrt{e}\simeq 0.606$: in words, the resolution threshold ensures the likelihood of observing any, two parallel hyperedges to be almost half of the probability of each, single hyperedge to be empty, hence suggesting that getting rid of parallel hyperedges is ‘easier’ than getting rid of empty ones.
Crossing the thresholds.
As Fig. 3 confirms, the probability for any, two hyperedges to be parallel decreases faster than the probability of each, single hyperedge to be empty: in words, while filling a large portion of the incidence matrix is not required to observe a limited amount of parallel hyperedges, when empty hyperedges are no longer observed, parallel hyperedges are no longer observed, too. More formally, $h_{r}^{\text{RHM}}=1/2<h_{f}^{\text{RHM}}=\ln L$: hence, progressively rising the density of 1s in the incidence matrix leads us crossing the resolution threshold before crossing the filling threshold. In more quantitative terms, evaluating $\langle N_{0}\rangle\xrightarrow[]{N\to+\infty}Le^{-h}$ in correspondence of $h_{r}^{\text{RHM}}=1/2$ returns the value $L/\sqrt{e}\simeq 606$ which is larger than $\langle d_{\alpha\beta}\rangle\xrightarrow[]{N\to+\infty}2h$ evaluated in correspondence of $h_{r}^{\text{RHM}}=1/2$ returning the value 1 by definition.
IV.1.2 2. The Hypergraph Configuration Model
@@
Let us, now, focus on the trends induced by our heterogeneous benchmarks. Following the procedure described in the previous sections, both the fitnesses of nodes and those of hyperedges were drawn from a Pareto distribution [^2] Other fat-tailed distributions were considered: qualitatively, results do not change. with $\alpha=2$. The right panel of Fig. 2 depicts the trend of $p_{1}$, as defined by Eq. (27). The right panel of Fig. 3 depicts the trend of $p_{0}$, as defined by Eq. (26), and the arithmetic mean $P_{0}\equiv 2\sum_{\alpha<\beta}P(d_{\alpha\beta}=0)/L(L-1)$ with $P(d_{\alpha\beta}=0)=\prod_{i=1}^{N}(1-q_{i}^{\alpha\beta})$, i.e. the average of the value $P(d_{\alpha\beta}=0)$ over all pairs of hyperedges - a choice dictated by the will of accounting for the heterogeneity of different pairs of hyperedges. Both quantities have been plotted as a function of $\langle T\rangle/NL\in[10^{-6},1]$ where $\langle T\rangle=\sum_{i=1}^{N}\sum_{\alpha=1}^{L}zf_{i}g_{\alpha}/(1+zf_{i}g_{\alpha})$. The dense (sparse) regime is recovered for large (small) values of $z$. Each dot of Figs. 2 and 3 represents an average taken over an ensemble of $1000$ configurations explicitly sampled from the HCM and is accompanied by the corresponding $95\%$ confidence interval. As in the case of the RHM, the agreement between the analytical trends (solid lines) and the numerical estimations (dots) confirms the correctness of our formulas.
Comparing the analytical estimation of (the trend of) the percentage of pairs of parallel hyperedges with its sample counterpart reveals the former to provide a quite accurate estimation of the latter. Besides, we explicitly notice that the trends characterising the HCM are sistematically shifted on the right: this is probably due to the presence of small fitnesses that, at least in case of Fig. 2, increase the probability of observing at least one, empty hyperedge, hence requiring a larger value of $z$ to let $p_{1}$ vanish. Such a result seems to be confirmed by the trends depicted in Fig. 3.
@@
Deriving an explicit expression for the filling threshold in the case of the HCM is a rather difficult task; still, we can proceed in a purely numerical fashion and individuate the value of the density of 1s in the incidence matrix guaranteeing that the minimum, expected hyperdegree amounts to 1: as evident from Fig. 3, such a threshold still lies in the right tail of the trends induced by the HCM - although its, precise, numerical value depends on those of the fitnesses.
IV.2 B. Solving the HCM on real-world hypergraphs
@@
In order to test our benchmarks on real-world configurations, we have focused on a number of data sets taken from Austin R. Benson’s website [^3] https://www.cs.cornell.edu/~arb/data/, i.e. the contact-primary-school, the email-Enron and the NDC-classes ones.
To have an explicit expression for the expected value of any quantity of interest, let us comment on the goodness of the approximation
@@
As Fig. 7 in Appendix A shows, it is is quite accurate for each data set considered here - in fact, one can safely assume that $x_{i}\simeq k_{i}^{*}/\sqrt{T^{*}}$, $\forall\>i$ and $h_{\alpha}^{*}/\sqrt{T^{*}}$, $\forall\>\alpha$.
IV.3 C. ‘Hypergraph to graph’ projection
@@
The canonical formalism that we have adopted leads to factorisable distributions, i.e. distributions that can be written as a product of pair-wise probability distributions; this allows the expectation of several quantities of interest to be evaluated analytically. To this aim, let us consider the matrix introduced in Battiston et al. (2020), reading
@@
with $\mathbf{K}$ being the diagonal matrix whose $i$-th entry reads $k_{i}$; according to the definition above
@@
counts the number of hyperedges nodes $i$ and $j$ are connected by - more explicitly, $w_{ij}=\sum_{\alpha=1}^{L}I_{i\alpha}I_{j\alpha}$, $i\neq j$ and $w_{ii}=\sum_{\alpha=1}^{L}I_{i\alpha}I_{i\alpha}-k_{i}=\sum_{\alpha=1}^{L}I_{i\alpha}-k_{i}=k_{i}-k_{i}=0$. The null models discussed so far can be employed to calculate $\langle w_{ij}\rangle$, $i\neq j$ that, in a perfectly general fashion, reads
@@
the total number of hyperedges shared by node $i$ with any other node in the hypergraph (in a sense, its ‘strength’ - see also Fig. 4) can be computed as
$\sigma_{i}=\sum_{\begin{subarray}{c}j=1\\ (j\neq i)\end{subarray}}^{N}w_{ij}=\sum_{\begin{subarray}{c}j=1\\ (j\neq i)\end{subarray}}^{N}\sum_{\alpha=1}^{L}I_{i\alpha}I_{j\alpha}$ (38)

### v3

@@ Results
Let us start by verifying the correctness of the estimations of the filling, multiple resolution and percolation thresholds provided by our benchmarks: to this aim, we have considered the values $N=300$ and $L=1000$.
The Random Hypergraph Model
@@
Each quantity has been plotted as a function of $p\in[10^{-6},1]$. The dense (sparse) regime is recovered for large (small) values of $p$. Each dot of Figs. 23 and 4 represents an average taken over an ensemble of $10^{3}$ configurations explicitly sampled from the RHM and is accompanied by the corresponding $95\%$ confidence interval, calculated via the bootstrap method [41].
The filling threshold
Panel a of Fig. 2 depicts the (analytical) trend of $\langle N_{\emptyset}\rangle/L=p_{\emptyset}=(1-p)^{N}$ (solid line): its agreement with the numerical estimations (dots) confirms the correctness of our formula.
@@
Although the value of the filling threshold has been determined by inspecting the asymptotic behaviour of $\langle N_{\emptyset}\rangle$, the quantity showing the neatest transition from the sparse to the dense regime is the probability of observing at least one empty hyperedge
${\color[rgb]{0,0,0}P(N_{\emptyset}>0)=1-P(N_{\emptyset}=0)=1-(1-p_{\emptyset})^{L}=1-[1-(1-p)^{N}]^{L}},$ (52)
@@
where we have exploited the fact that $P(N_{\emptyset}>0)$ is nothing but the complementary of the probability that no hyperedge is empty. Since $p_{\emptyset}\xrightarrow[]{N\to+\infty}e^{-h}$, evaluating such an expression in correspondence of $h_{f}^{\text{RHM}}=\ln L\simeq 6.907$ returns the value $1/L=10^{-3}$ (see panel a of Fig. 2). As a consequence, the result $P(N_{\emptyset}>0)\xrightarrow[]{N\to+\infty}1-(1-e^{-h})^{L}$ is numerically recovered (see panel b of Fig. 2). In words, although the filling threshold ensures that each, single hyperedge is empty with an overall small probability, the likelihood of observing at least one, empty hyperedge is still large (i.e. 2/3$\simeq$ 2/3)≃ 2 / 3 ): the steepness of the trend of $P(N_{\emptyset}>0)$, however, suggests it to quickly vanish as the density of $1$s in the incidence matrix crosses the value $p_{f}^{\text{RHM}}=h_{f}^{\text{RHM}}/N=\ln L/N\simeq 0.023$.
Figure 2: Impact of empty hyperedges on the RHM ensemble. More in detail, the trends of $\langle N_{\emptyset}\rangle/L=p_{\emptyset}=(1-p)^{N}\xrightarrow[]{N\to+\infty}e^{-h}$, i.e. the probability for the generic hyperedge to be emtpy is represented in panel a and the one of $P(N_{\emptyset}>0)=1-[1-(1-p)^{N}]^{L}\xrightarrow[]{N\to+\infty}1-(1-e^{-h})^{L}$, i.e. the probability of observing at least one empty hyperedge is represented in panel b. Evaluating them in correspondence of $p_{f}^{\text{RHM}}=h_{f}^{\text{RHM}}/N=\ln L/N\simeq 0.023$ (vertical line) returns, respectively, the values $1/L=10^{-3}$ and $1-(1-1/L)^{L}\simeq 0.6323$. The dense (sparse) regime is recovered for large (small) values of $p$. Each dot represents an average taken over an ensemble of $10^{3}$ configurations (explicitly sampled from the RHM) and is accompanied by the corresponding $95\%$ confidence interval.
The multiple resolution threshold
Panel a of Fig. 3 depicts the trend of $2\langle N_{\mathbin{/\!/}}\rangle/L(L-1)=[1-2p(1-p)]^{N}=p_{\mathbin{/\!/}}$ (solid line): again, its agreement with the numerical estimations (dots) confirms the correctness of our formula.
@@
Evaluating $p_{\mathbin{/\!/}}\xrightarrow[]{N\to+\infty}e^{-2h}$ in correspondence of $h_{m}^{\text{RHM}}\lesssim\ln L\simeq 6.907$ returns the value $1/L^{2}=10^{-6}$ (see panel a of Fig. 3): in words, the aforementioned, critical value causes the likelihood of observing any two parallel hyperedges to be almost the square of the probability for each single hyperedge to be empty (i.e. $1/L$, see the comments under Eq. 17).
As the overlaps between pairs of hyperedges cannot be treated as i.i.d. random variables, evaluating the probability of observing at least one pair of parallel hyperedges forces us to proceed in a purely numerical fashion. Calculating $P(N_{\mathbin{/\!/}}>0)$ in correspondence of $p_{m}^{\text{RHM}}=h_{m}^{\text{RHM}}/N\lesssim\ln L/N\simeq 0.023$ returns the value $0.6$ (see the panel b of Fig. 3).
Carrying out such an estimation in the aforementioned regime is, however, instructive as it leads to the expression $P(N_{\mathbin{/\!/}}>0)=1-(1-p_{\mathbin{/\!/}})^{L(L-1)/2}=1-\{1-[1-2p(1-p)]^{N}\}^{L(L-1)/2}\xrightarrow[]{N\to+\infty}1-[1-e^{-2h}]^{L(L-1)/2}$ that, evaluated in correspondence of $h_{m}^{\text{RHM}}\lesssim\ln L\simeq 6.907$, returns the value $P(N_{\mathbin{/\!/}}>0)=1-(1-1/L^{2})^{L(L-1)/2}\simeq 0.393$ whose difference with $0.6$, obtained numerically, lets us fully appreciate the role played by correlations.
Figure 3: Impact of parallel hyperedges on the RHM ensemble. More in detail, the trends of $2\langle N_{\mathbin{/\!/}}\rangle/L(L-1)=[1-2p(1-p)]^{N}=p_{\mathbin{/\!/}}\xrightarrow[]{N\to+\infty}e^{-2h}$, i.e. the probability for the generic pair of hyperedges to be parallel and $P(N_{\mathbin{/\!/}}>0)$, i.e. the probability of observing at least one pair of parallel hyperedges, as functions of $p$, are represented, respectively, in panels a and b. Evaluating them in correspondence of $p_{m}^{\text{RHM}}=h_{m}^{\text{RHM}}/N\lesssim\ln L/N\simeq 0.023$ (vertical line) returns, respectively, the values $1/L^{2}=10^{-6}$ and $\simeq 0.6$. The dense (sparse) regime is recovered for large (small) values of $p$. Each dot represents an average taken over an ensemble of $10^{3}$ configurations (explicitly sampled from the RHM) and is accompanied by the corresponding $95\%$ confidence interval.
The percolation threshold
Let us, now, focus on the projection of our hypergraph onto the layer of hyperedges. The generic hyperedge is isolated either because is not ‘connected’ with any node or because is a singleton (i.e. it is ‘connected’ with a node with which no other hyperedge is ‘connected’): in symbols,
${\color[rgb]{0,0,0}p_{0}\equiv\{1-p[1-(1-p)^{L-1}]\}^{N}=[(1-p)+p(1-p)^{L-1}]^{N}}.$ (53)
In order to evaluate $p_{0}$ in correspondence of the percolation threshold, let us consider that $L=s^{2}N$, with $s^{2}=10/3$; one, then, finds
${\color[rgb]{0,0,0}\lim_{N\to+\infty}p_{0}=\lim_{N\to+\infty}\left[\left(1-\frac{1}{Ns}\right)+\frac{1}{Ns}\left(1-\frac{1}{Ns}\right)^{s^{2}N-1}\right]^{N}=e^{-\frac{1-e^{-s}}{s}}},$ (54)
whose numerical value amounts to $0.631$ (see panel a of Fig. 4): in words, the value $p_{0}(p=p_{p}^{\text{RHM}})\lesssim 2/3$ implies that the expected number of isolated hyperedges in the projection $\langle N_{0}\rangle=Lp_{0}$ tends to $32L/$ 3 as $p$ tends to $p_{p}^{\text{RHM}}$.
Figure 4: Impact of isolated hyperedges on the RHM ensemble. More in detail, trends of $\langle N_{0}\rangle/L=p_{0}$, i.e. the probability for the generic hyperedge to be isolated in the projection, $P(N_{0}>0)$, i.e. the probability of observing at least one, isolated hyperedge in the projection and $|\text{LCC}|/N$, i.e. the percentage of nodes belonging to the largest connected component (LCC), are represented as functions of $p$, respectively in panels a, b and c. Evaluating the first two in correspondence of $p_{p}^{\text{RHM}}=h_{p}^{\text{RHM}}/N=1/\sqrt{NL}\simeq 0.002$ (vertical line) returns, respectively, the values $e^{(e^{-s}-1)/s}\simeq 0.631$ and $\simeq 1$. The dense (sparse) regime is recovered for large (small) values of $p$. Each dot represents an average taken over an ensemble of $10^{3}$ configurations (explicitly sampled from the RHM) and is accompanied by the corresponding $95\%$ confidence interval.
Pairs of hyperedges cannot be treated as independent. Let us, in fact, consider the bipartite representation of a hypergraph: as adjacent pairs of hyperedges (say $\alpha$, $\beta$ and $\beta$, $\gamma$) may share some neighbours (on the opposite layer), the number of common neighbours of $\alpha$ and $\beta$ will, in general, covariate with the number of common neighbours of $\beta$ and $\gamma$; therefore, evaluating the probability of observing at least one, isolated hyperedge in the projection forces us to proceed in a purely numerical fashion. Calculating $P(N_{0}>0)$ in correspondence of $p_{p}^{\text{RHM}}=h_{p}^{\text{RHM}}/N=1/\sqrt{NL}\simeq 0.002$ practically returns $1$ (see panel b of Fig. 4). From the perspective of a hypergraph connectedness, the percolation threshold is ‘less strict’ than the filling threshold, allowing for a larger number of disconnected nodes ($2/3$ of the total versus $1$).
As before, estimating the percolation threshold in the regime where the pairs of hyperedges behave as i.i.d. Binomial random variables is instructive. In this case, projecting a bipartite networks onto the layer of hyperedges amounts to connect any two of them with probability $1-(1-p^{2})^{N}$ i.e. the complementary of the probability $(1-p^{2})^{N}$ of not sharing any node. The number of isolated nodes, thus, obeys the relationship $N_{0}\sim\text{Bin}(L,p_{0})$, with
${\color[rgb]{0,0,0}p_{0}\equiv[(1-p^{2})^{N}]^{L}}$ (55)
being the probability for the generic hyperedge to be isolated: in words, such an expression returns the probability for the generic hyperedge to not share any node - with probability $(1-p^{2})^{N}$ with any other hyperedge - with a probability amounting to the previous one raised to the power of $L$. As a consequence, evaluating $p_{0}$ in correspondence of the percolation threshold returns a value tending to $1/3$, a result further implying that the expected number of isolated hyperedges in the projection $\langle N_{0}\rangle=Lp_{0}$ tends to the value $L/3$ both letting us fully appreciate the role played by correlations.
Within such a context, the probability of observing at least one isolated hyperedge in the projection satisfies the chain of relationships $P(N_{0}>0)=1-P(N_{0}=0)=1-(1-p_{0})^{L}=1-[1-(1-p^{2})^{NL}]^{L}\xrightarrow[]{N\to+\infty}1-(1-e^{-1})^{L}$: as the last expression quickly converges to 1 for large values of $L$, the same qualitative behaviour observed before is thus recovered.
Finally, one may wonder which kind of mesoscale structure is identified by the percolation threshold: the answer is provided by panel c of Fig. 4, showing the appearance of a large connected component - as also pointed out in [12], the presence of a large connected component is inspected by projecting our hypergraph onto the layer of nodes, whose connectedness is ensured by requiring the connectedness of hyperedges.
The role of thresholds in a hypergraph evolution
Let us, now, make a couple of observations. The first one concerns the result according to which $h_{m}^{\text{RHM}}\lesssim h_{f}^{\text{RHM}}$: such a relationship suggests that, while filling a large portion of the incidence matrix is not required to observe a limited amount of parallel hyperedges (see panel b of Fig. 3), when empty hyperedges are no longer observed, parallel hyperedges are no longer observed as well.
The second one concerns the result according to which either $h_{f}^{\text{RHM}}\leq h_{p}^{\text{RHM}}$ or $h_{f}^{\text{RHM}}\geq h_{p}^{\text{RHM}}$. By progressively rising the parameter $p$, two, different thresholds are, thus, met: if $h_{f}^{\text{RHM}}\leq h_{p}^{\text{RHM}}$, the filling threshold $p_{f}^{\text{RHM}}=\ln L/N$ is met before the percolation threshold $p_{p}^{\text{RHM}}=1/\sqrt{NL}$, i.e. hyperedges are filled before they start sharing nodes - as a consequence, singletons appear; if $h_{f}^{\text{RHM}}\geq h_{p}^{\text{RHM}}$, the percolation threshold $p_{p}^{\text{RHM}}=1/\sqrt{NL}$ is met before the filling threshold $p_{f}^{\text{RHM}}=\ln L/N$, i.e. hyperedges start sharing nodes before they are filled - as a consequence, no singleton appears before the filling threshold is crossed.
Notice that, for simple graphs, $k_{p}^{\text{RGM}}=1\leq k_{c}^{\text{RGM}}=\ln N$, i.e. by progressively rising the parameter $q$, the percolation threshold is always met before the connectivity threshold.
The Hypergraph Configuration Model
@@
In order to carry out the numerical simulations in the case of the HCM, we have followed the procedure described in the previous sections and drawn both the fitnesses of nodes and those of hyperedges from a Pareto distribution with $\alpha=2$ other fat-tailed distributions were considered: qualitatively, results do not change. Each quantity has been plotted as a function of $\rho(z)=\langle T\rangle/NL\in[10^{-6},1]$ where $\langle T\rangle=\sum_{i=1}^{N}\sum_{\alpha=1}^{L}zf_{i}g_{\alpha}/(1+zf_{i}g_{\alpha})$ varies with $z$. The dense (sparse) regime is recovered for large (small) values of $z$. Each dot of Figs. 56 and 7 represents an average taken over an ensemble of $10^{3}$ configurations explicitly sampled from the HCM and is accompanied by the corresponding $95\%$ confidence interval.
The filling threshold
@@
Panel a of Fig. 5 depicts the (analytical) trend of $\langle N_{\emptyset}\rangle/L=\sum_{\alpha=1}^{L}p_{\emptyset}^{\alpha}/L=\overline{p_{\emptyset}}$ (solid line): as in the case of the RHM, its agreement with the numerical estimations (dots) confirms the correctness of our formula. Deriving an explicit expression for the filling threshold in the case of the HCM is a rather difficult task; still, we can proceed in a purely numerical fashion and individuate the value of the density of $1$s in the incidence matrix guaranteeing that the expected number of empty hyperedges (divided by $L$) amounts to $1$ (divided by $L$): although its, precise, numerical value depends on the values of the fitnesses, such a threshold still lies in the right tail of the trend induced by the HCM and reads $p_{f}^{\text{HCM}}\simeq 0.032$ (see panel a of Fig. 5).
Figure 5: Impact of empty hyperedges on the HCM ensemble. More in detail, trends of $\langle N_{\emptyset}\rangle/L=\overline{p_{\emptyset}}$, i.e. the probability for the generic hyperedge to be empty and $P(N_{\emptyset}>0)\simeq 1-\prod_{\alpha=1}^{L}(1-e^{-h_{\alpha}})$, i.e. the probability of observing at least one empty hyperedge, are represented as functions of the connectance $\rho$, respectively in panels a and b. Evaluating the latter in correspondence of the filling threshold, reading $p_{f}^{\text{HCM}}\simeq 0.032$ (vertical line), returns the value $\simeq 0.642$. The dense (sparse) regime is recovered for large (small) values of $z$. Each dot represents an average taken over an ensemble of $10^{3}$ configurations (explicitly sampled from the HCM) and is accompanied by the corresponding $95\%$ confidence interval.
Even in this case, the quantity showing the neatest transition from the sparse to the dense regime is the probability of observing at least one empty hyperedge
${\color[rgb]{0,0,0}P(N_{\emptyset}>0)=1-P(N_{\emptyset}=0)=1-\prod_{\alpha=1}^{L}(1-p_{\emptyset}^{\alpha})=1-\prod_{\alpha=1}^{L}\left[1-\prod_{i=1}^{N}(1-p_{i\alpha})\right]}$ (56)
that, in the sparse regime, can be approximated as $P(N_{\emptyset}>0)\simeq 1-\prod_{\alpha=1}^{L}(1-e^{-h_{\alpha}})$: as panel b of Fig. 5 shows, evaluating $P(N_{\emptyset}>0)$ in correspondence of the filling threshold returns $0.642$. Finally, let us explicitly notice that the value of the filling threshold is shifted on the right with respect to its homogeneous counterpart, an evidence probably due to the presence of small fitnesses that increase the probability of observing at least one, empty hyperedge, hence requiring a larger value of $z$ to let $P(N_{\emptyset}>0)$ vanish.
Figure 6: Impact of parallel hyperedges on the HCM ensemble. More in detail, trends of $2\langle N_{\mathbin{/\!/}}\rangle/L(L-1)=\overline{p_{\mathbin{/\!/}}}$, i.e. the probability for the generic pair of hyperedges to be parallel and $P(N_{\mathbin{/\!/}}>0)$, i.e. the probability of observing at least one pair of parallel hyperedges, are represented as functions of the connectance $\rho$, respectively in panels a and b. Evaluating the latter in correspondence of the multiple resolution threshold, reading $p_{m}^{\text{HCM}}\simeq 0.031$ (vertical line), returns the $0.493$. The dense (sparse) regime is recovered for large (small) values of $z$. Each dot represents an average taken over an ensemble of $10^{3}$ configurations (explicitly sampled from the HCM) and is accompanied by the corresponding $95\%$ confidence interval.
The multiple resolution threshold
Panel a of Fig. 6 depicts the (analytical) trend of $2\langle N_{\mathbin{/\!/}}\rangle/L(L-1)=2\sum_{\alpha=1}^{L}\sum_{\begin{subarray}{c}\beta=1\\ \beta>\alpha\end{subarray}}^{L}p_{\mathbin{/\!/}}^{\alpha\beta}/L(L-1)=\overline{p_{\mathbin{/\!/}}}$ (solid line): as in the case of the RHM, its agreement with the numerical estimations (dots) confirms the correctness of our formula.
Adopting the strategy described in the previous paragraph, i.e. that of requiring that the expected number of parallel hyperedges (divided by $L(L-1)/2$) amounts to $1$ (divided by $L(L-1)/2$), we found that $p_{m}^{\text{HCM}}\simeq 0.031$ (see panel a of Fig. 6). Numerically clculating $P(N_{\mathbin{/\!/}}>0)$ in correspondence of the multiple resolution threshold returns the value $0.493$ (see panel b of Fig. 6).
Let us notice that the value of the multiple resolution threshold no longer coincides with the value of the filling threshold, although it is still shifted on the right with respect to its homogeneous counterpart.
Figure 7: Impact of isolated hyperedges on the HCM ensemble. More in detail, trends of $\langle N_{0}\rangle/L=\overline{p_{0}}$, i.e. the probability for the generic hyperedge to be isolated in the projection, $P(N_{0}>0)$, i.e. the probability of observing at least one, isolated hyperedge in the projection, and $|\text{LCC}|/N$, i.e. the percentage of nodes belonging to the largest connected component (LCC), are repesented as functions of the connectance $\rho$, respectively in panels a, b and c. Evaluating the first two in correspondence of $p_{p}^{\text{HCM}}\simeq 0.001$ (vertical line), returns the values $\simeq 0.766$ and $\simeq 1$. The dense (sparse) regime is recovered for large (small) values of $p$. Each dot represents an average taken over an ensemble of $10^{3}$ configurations (explicitly sampled from the HCM) and is accompanied by the corresponding $95\%$ confidence interval.
The percolation threshold
The probability for the generic hyperedge to be isolated in the projection, now, reads
${\color[rgb]{0,0,0}p_{0}^{\alpha}\equiv\prod_{i=1}^{N}\left\{1-p_{i\alpha}\left[1-\prod_{\begin{subarray}{c}\beta=1\\ \beta\neq\alpha\end{subarray}}^{L}(1-p_{i\beta})\right]\right\}=\prod_{i=1}^{N}\left\{(1-p_{i\alpha})+p_{i\alpha}\prod_{\begin{subarray}{c}\beta=1\\ \beta\neq\alpha\end{subarray}}^{L}(1-p_{i\beta})\right\}}$ (57)
that, in the sparse regime, can be approximated as $p_{0}^{\alpha}\simeq\prod_{i=1}^{N}[(1-p_{i\alpha})+p_{i\alpha}e^{-k_{i}}]$. Adopting the strategy described in the previous paragraphs, i.e. that of requiring that the expected value of the total number of nodes shared by any hyperedge with any other hyperedge amounts to $1$ in symbols, $\overline{\langle\sigma\rangle}=\sum_{\alpha=1}^{L}\langle\sigma_{\alpha}\rangle/L=1$ we found that $p_{p}^{\text{HCM}}\simeq 0.001$: since $\langle N_{0}\rangle=\sum_{\alpha=1}^{L}p_{0}^{\alpha}$, evaluating $\langle N_{0}\rangle/L=\sum_{\alpha=1}^{L}p_{0}^{\alpha}/L=\overline{p_{0}}$ and $P(N_{0}>0)$ in correspondence of the percolation threshold returns respectively the values $0.766$ (see panel a of Fig. 7) and $\simeq 1$ (see panel b of Fig. 7). For what concerns the hypergraph connectedness, the same conclusion drawn in the case of the RHM holds true, as the percolation threshold allows for a larger number of disconnected nodes ($3/4$ of the total versus $1$).
Analogously, the mesoscale structure individuated by the percolation threshold consists of a large connected component constituted (see panel c of Fig. 7).
Solving the HCM on real-world hypergraphs
@@
In order to test our benchmarks on real-world configurations, we have focused on a number of data sets taken from Austin R. Benson’s website (https://www.cs.cornell.edu/~arb/data/), i.e. the contact-primary-school, the email-Enron and the NDC-classes ones.
Although the parameters defining the HCM must be numerically determined by solving the system of equations induced by the likelihood maximisation, when the system under analysis is sparse they can be approximated as described in A of the Appendix. Such an approximation leads to the expression
$p_{i\alpha}\simeq x_{i}y_{\alpha}=\frac{k_{i}^{*}h_{\alpha}^{*}}{T^{*}},\>\forall\>i,\alpha$ (58)
@@
that, as Fig. 11 in the Appendix shows, is quite accurate for each data set considered here - in fact, one can safely assume that $x_{i}\simeq k_{i}^{*}/\sqrt{T^{*}}$, $\forall\>i$ and $h_{\alpha}^{*}/\sqrt{T^{*}}$, $\forall\>\alpha$.
‘Hypergraph to graph’ projection
@@
The canonical formalism that we have adopted leads to factorisable distributions, i.e. distributions that can be written as a product of pair-wise probability distributions; this allows the expectation of several quantities of interest to be evaluated analytically.
Let us start by considering the matrix, introduced in [4], reading
@@
with $\mathbf{K}$ being the diagonal matrix whose $i$-th entry reads $k_{i}$; according to the definition above, it induces a projection of a hypergraph onto a weighted graph, whose generic entry
@@
returns the number of hyperedges both $i$ and $j$ belong to - more explicitly, $w_{ij}=\sum_{\alpha=1}^{L}I_{i\alpha}I_{j\alpha}$, $i\neq j$ and $w_{ii}=\sum_{\alpha=1}^{L}I_{i\alpha}I_{i\alpha}-k_{i}=\sum_{\alpha=1}^{L}I_{i\alpha}-k_{i}=k_{i}-k_{i}=0$. In other words, $\mathbf{W}$ represents the object most closely resembling a traditional adjacency matrix. The null models discussed so far can be employed to calculate $\langle w_{ij}\rangle$, $i\neq j$ that, in a perfectly general fashion, reads
@@
as we said, the total number of hyperedges shared by node $i$ with any other node in the hypergraph (in a sense, its ‘strength’ - see also Fig. 8) can be computed as
$\sigma_{i}=\sum_{\begin{subarray}{c}j=1\\ j\neq i\end{subarray}}^{N}w_{ij}=\sum_{\begin{subarray}{c}j=1\\ j\neq i\end{subarray}}^{N}\sum_{\alpha=1}^{L}I_{i\alpha}I_{j\alpha}$ (62)